10) Given these two individual demand curves of consumers A and B, what is the equation for the market demand curve? i. P = 100 – 10qA ii. P = 100 - (5/4)qB

A. P = 90 – 10/9Q

B. P = 200 – 11.25Q

C. P = 100 – 11.25Q

D. P = 90 – 0.9Q

E. None of the above
Answer #1

Market demand curve is obtained by horizontally adding individual demand curve.In this case the intercept (100 ) would remain the same but the slope will change.

New slope=10+5/4=11.25

New market demand=100-11.25Q

Answer-C